# Detecting Visual Information Manipulation Attacks in Augmented Reality: A Multimodal Semantic Reasoning Approach

### v1

@@ 7 Limitations and Future Work
While VIM-Sense demonstrates strong performance in detecting visual information manipulation attacks in AR, several limitations remain. First, the system relies on VLMs operating in a zero-shot setting, which introduces inherent variability in responses. Although our use of structured chain-of-thought prompting improves stability, generative models may still yield inconsistent judgments, particularly in ambiguous cases. In response to this challenge, we will explore fine-tuning large-scale VLMs on AR-VIM or related datasets to improve their robustness in AR-specific semantic reasoning, especially as recent studies have demonstrated that even commercial models can be effectively fine-tuned using small-scale datasets [30, 10]. Second, although VIM-Sense uses a lightweight edge-cloud architecture, the latency remains in the range of several seconds. This is suitable for general AR use cases, especially those that only require periodic attack detection and do not involve rapid changes in scene information and virtual content, such as indoor navigation, assembly training, and visual searching, but may be inadequate for high-speed or real-time use cases such as AR-assisted driving or human-robot collaboration. We will monitor the advances in the generative AI field, and explore techniques such as model compression and knowledge distillation, particularly for open-source VLMs, to integrate more lightweight alternatives when commercial APIs are impractical. These efforts will continuously incorporate faster and more efficient models into our system.
@@
Meanwhile, the attack taxonomy of AR-VIM was mainly focused on the semantic information change, which can be further enriched by involving other types of attacks in AR, such as cognitive overload[17, 11], which introduce excessive or cluttered virtual elements and cause users to overlook critical real-world cues. Our future work will also consider risks introduced by non-adversarial AR content, where benign overlays may unintentionally trigger the same mechanisms of information obfuscation or introduce extra wrong information. Such content can lead to similar detrimental effects on user perception and decision-making, even though they are not intended to be malicious. Some samples in AR-VIM reflected this kind of effect yet a more systematic discussion is still required. Also, AR-VIM does not yet cover multi-modal attacks that may involve additional sensory channels such as audio[60, 59] or haptics[7, 8], which are becoming increasingly common in real-world AR applications and may jointly influence user perception. Thus, we will explore multimodal cues to create more attack cases, and extend the detection framework to handle such cross-modal attacks.
Finally, while VIM-Sense currently relies heavily on VLMs as the core reasoning engine, we will move beyond prompt engineering by incorporating other lightweight interpretable modules to assist the VLM. This will also enable VIM-Sense to produce more informative outputs, such as identifying the specific spatial regions affected by the attack, rather than providing only a binary decision. One direction is to integrate multi-modal object detection models for visual grounding[46, 39], enabling the system to verify whether the VLM’s reasoning corresponds to actual content in the AR scene. We will also introduce a form of context memory that tracks scene context or prior detections to support more consistent and context-aware prompt construction[34]. These extensions will make the reasoning process more robust, reduce reliance on independent prompts, and improve interpretability in real-world AR deployments.
@@ 8 Conclusion
In this work, we proposed a multimodal semantic reasoning approach to address the challenge of detecting visual information manipulation attacks in AR environments. We first proposed a mathematical formulation for such attacks, defining them based on attack formats and attack purposes. Building on this framework, we introduced the AR-VIM dataset, which consists of 452 raw-AR video pairs across 202 scenes, covers the defined attack types, and is validated through an IRB-approved user-based data labeling task. To detect the attacks, we developed VIM-Sense, which leverages VLMs and OCR-based prompt construction within an edge-cloud architecture. Our experiments show that VIM-Sense largely outperforms baseline methods in attack detection accuracy and achieves a moderate responsiveness. Finally, VIM-Sense is also deployed and tested with a real-world mobile AR application, achieving a moderate latency and demonstrating potential for safeguarding semantic integrity in AR applications.
@@ Acknowledgements.
We thank our user study participants for their invaluable assistance in this research. This work was supported in part by NSF grants CSR-2312760, CNS-2112562, and IIS-2231975, NSF CAREER Award IIS-2046072, NSF NAIAD Award 2332744, a CISCO Research Award, a Meta Research Award, Defense Advanced Research Projects Agency Young Faculty Award HR0011-24-1-0001, and the Army Research Laboratory under Cooperative Agreement Number W911NF-23-2-0224. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Defense Advanced Research Projects Agency, the Army Research Laboratory, or the U.S. Government. This paper has been approved for public release; distribution is unlimited. No official endorsement should be inferred. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein.References
@@
- [11] J. Buchner, K. Buntins, and M. Kerres. The impact of augmented reality on cognitive load and performance: A systematic review. Journal of Computer Assisted Learning, 38(1):285–303, 2022.
@@
- [14] J. Chung, S. Hyun, and J.-P. Heo. Style injection in diffusion: A training-free approach for adapting large-scale diffusion models for style transfer. In Proceedings of the IEEE/CVF conference on computer vision and pattern recognition, pp. 8795–8805, 2024.
@@
- [17] J. T. Doswell and A. Skinner. Augmenting human cognition with adaptive augmented reality. In International Conference on Augmented Cognition, pp. 104–113. Springer, 2014.
@@
- [22] M. Golyadkin, S. Saraev, and I. Makarov. Benchmarking and data synthesis for colorization of manga sequential pages for augmented reality. In 2024 IEEE International Symposium on Mixed and Augmented Reality Adjunct (ISMAR-Adjunct), pp. 608–611, 2024.
@@
- [29] P. Krishnan, R. Kovvuri, G. Pang, B. Vassilev, and T. Hassner. TextStyleBrush: transfer of text aesthetics from a single example. IEEE Transactions on Pattern Analysis and Machine Intelligence, 45(7):9122–9134, 2023.
@@
- [36] X. Li, Q. Zhang, D. Kang, W. Cheng, Y. Gao, J. Zhang, Z. Liang, J. Liao, Y.-P. Cao, and Y. Shan. Advances in 3D generation: A survey. arXiv preprint arXiv:2401.17807, 2024.
@@
- [39] S. Liu, Z. Zeng, T. Ren, F. Li, H. Zhang, J. Yang, Q. Jiang, C. Li, J. Yang, H. Su, et al. Grounding DINO: Marrying DINO with grounded pre-training for open-set object detection. In Proceedings of European Conference on Computer Vision. Springer, 2024.
@@
- [45] A. Sharma, L. Yoffe, and T. Höllerer. OCTO+: A Suite for Automatic Open-Vocabulary Object Placement in Mixed Reality . In IEEE International Conference on Artificial Intelligence and eXtended and Virtual Reality (AIxVR), pp. 157–165, 2024.
- [46] H. Shen, T. Zhao, M. Zhu, and J. Yin. GroundVLP: Harnessing zero-shot visual grounding from vision-language pre-training and open-vocabulary object detection. In Proceedings of the AAAI Conference on Artificial Intelligence, vol. 38, 2024.
@@
- [48] S. Srinidhi, E. Lu, and A. Rowe. XaiR: An XR platform that integrates large language models with the physical world. In IEEE International Symposium on Mixed and Augmented Reality (ISMAR), pp. 759–767, 2024.
- [49] D. Tochilkin, D. Pankratz, Z. Liu, Z. Huang, A. Letts, Y. Li, D. Liang, C. Laforte, V. Jampani, and Y.-P. Cao. TripoSR: Fast 3D object reconstruction from a single image. arXiv preprint arXiv:2403.02151, 2024.
- [50] W. Tseng, E. Bonnail, M. McGill, M. Khamis, E. Lecolinet, S. Huron, and J. Gugenheimer. The dark side of perceptual manipulations in virtual reality. In Proceedings of the ACM CHI Conference on Human Factors in Computing Systems, article no. 612, 2022.
@@
- [56] J. Xiang, Z. Lv, S. Xu, Y. Deng, R. Wang, B. Zhang, D. Chen, X. Tong, and J. Yang. Structured 3D latents for scalable and versatile 3d generation. In Proceedings of the IEEE/CVF Computer Vision and Pattern Recognition Conference, 2025.

### v3

@@ 6 Limitations and Future Work
While VIM-Sense demonstrates strong performance in detecting visual information manipulation attacks in AR, several limitations remain. First, the system relies on VLMs operating in a zero-shot setting, which introduces inherent variability in responses. Although our use of structured chain-of-thought prompting improves stability, generative models may still yield inconsistent judgments, particularly in ambiguous cases. In response to this challenge, we will explore fine-tuning large-scale VLMs on AR-VIM or related datasets to improve their robustness in AR-specific semantic reasoning, especially as recent studies have demonstrated that even commercial models can be effectively fine-tuned using small-scale datasets [31, 10]. Second, although VIM-Sense uses a lightweight edge-cloud architecture, the latency remains in the range of several seconds. This is suitable for general AR use cases, especially those that only require periodic attack detection and do not involve rapid changes in scene information and virtual content, such as indoor navigation, assembly training, and visual searching, but may be inadequate for high-speed or real-time use cases such as AR-assisted driving or human-robot collaboration. We will monitor the advances in the generative AI field and explore techniques such as model compression and knowledge distillation, particularly for open-source VLMs, to integrate more lightweight alternatives when commercial APIs are impractical. These efforts will continuously incorporate faster and more efficient models into our system.
@@
Meanwhile, the attack taxonomy of AR-VIM was mainly focused on the semantic information change, which can be further enriched by involving other types of attacks in AR. For example, cognitive overload[17, 11], which introduces excessive or cluttered virtual elements and cause users to overlook critical real-world cues. Our future work will also consider risks introduced by non-adversarial AR content, where benign overlays may unintentionally trigger the same mechanisms of information obfuscation or introduce extra wrong information. Such content can lead to similar detrimental effects on user perception and decision-making, even though it is not intended to be malicious. Some samples in AR-VIM reflected this kind of effect, yet a more systematic data collection is still required. Also, AR-VIM does not yet cover multi-modal attacks that may involve additional sensory channels such as audio[61, 60] or haptics[7, 8], which are becoming increasingly common in real-world AR applications and may jointly influence user perception. Thus, we will explore multimodal cues to create more attack cases and extend the detection framework to handle such cross-modal attacks.
Finally, while VIM-Sense currently relies heavily on VLMs as the core reasoning engine, we will move beyond prompt engineering by incorporating other lightweight interpretable modules to assist the VLM. This will also enable VIM-Sense to produce more informative outputs, such as identifying the specific spatial regions affected by the attack, rather than providing only a binary decision. One direction is to integrate multi-modal object detection models for visual grounding[47, 40], enabling the system to verify whether the VLM’s reasoning corresponds to actual content in the AR scene. We will also introduce a form of context memory that tracks scene context or prior detections to support more consistent and context-aware prompt construction[35]. These extensions will make the reasoning process more robust, reduce reliance on individual prompts, and improve interpretability in real-world AR deployments.
@@ 7 Conclusion
In this work, we proposed a multimodal semantic reasoning approach to address the challenge of detecting visual information manipulation attacks in AR environments. We first proposed a mathematical formulation for such attacks, defining them based on attack formats and attack purposes. Building on this framework, we introduced the AR-VIM dataset, which consists of 452 raw-AR video pairs across 202 scenes, covers the defined attack types, and is validated through an IRB-approved user-based data labeling task. To detect the attacks, we developed VIM-Sense, which leverages VLMs and OCR-based prompt construction within an edge-cloud architecture. Our experiments show that VIM-Sense largely outperforms baseline methods in attack detection accuracy and achieves a moderate responsiveness. Finally, VIM-Sense is also deployed and tested with a real-world mobile AR application, demonstrating potential for safeguarding semantic integrity in AR applications.
@@ Acknowledgements.
We thank the participants of our user-based label validation for their invaluable effort and assistance in this research. This work was supported in part by NSF grants CSR-2312760, CNS-2112562, and IIS-2231975, NSF CAREER Award IIS-2046072, NSF NAIAD Award 2332744, a Cisco Research Award, a Meta Research Award, Defense Advanced Research Projects Agency Young Faculty Award HR0011-24-1-0001, and the Army Research Laboratory under Cooperative Agreement Number W911NF-23-2-0224. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Defense Advanced Research Projects Agency, the Army Research Laboratory, or the U.S. Government. This paper has been approved for public release; distribution is unlimited. No official endorsement should be inferred. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein.References
@@
- [11] J. Buchner, K. Buntins, and M. Kerres. The impact of augmented reality on cognitive load and performance: A systematic review. Journal of Computer Assisted Learning, 38(1), 2022.
@@
- [14] J. Chung, S. Hyun, and J.-P. Heo. Style injection in diffusion: A training-free approach for adapting large-scale diffusion models for style transfer. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR), 2024.
@@
- [17] J. T. Doswell and A. Skinner. Augmenting human cognition with adaptive augmented reality. In International Conference on Augmented Cognition, 2014.
@@
- [22] M. Golyadkin, S. Saraev, and I. Makarov. Benchmarking and data synthesis for colorization of manga sequential pages for augmented reality. In 2024 IEEE International Symposium on Mixed and Augmented Reality Adjunct (ISMAR-Adjunct), 2024.
@@
- [25] W. Huang, H. Liu, M. Guo, and N. Gong. Visual hallucinations of multi-modal large language models. In Findings of the Association for Computational Linguistics ACL, 2024.
@@
- [30] P. Krishnan, R. Kovvuri, G. Pang, B. Vassilev, and T. Hassner. TextStyleBrush: transfer of text aesthetics from a single example. IEEE Transactions on Pattern Analysis and Machine Intelligence, 45(7), 2023.
@@
- [37] X. Li, Q. Zhang, D. Kang, W. Cheng, Y. Gao, J. Zhang, Z. Liang, J. Liao, Y. Cao, and Y. Shan. Advances in 3D generation: A survey. arXiv preprint arXiv:2401.17807, 2024.
@@
- [40] S. Liu, Z. Zeng, T. Ren, F. Li, H. Zhang, J. Yang, Q. Jiang, C. Li, J. Yang, H. Su, et al. Grounding DINO: Marrying DINO with grounded pre-training for open-set object detection. In Proceedings of European Conference on Computer Vision, 2024.
@@
- [46] A. Sharma, L. Yoffe, and T. Höllerer. OCTO+: A Suite for Automatic Open-Vocabulary Object Placement in Mixed Reality . In IEEE International Conference on Artificial Intelligence and eXtended and Virtual Reality (AIxVR), 2024.
- [47] H. Shen, T. Zhao, M. Zhu, and J. Yin. GroundVLP: Harnessing zero-shot visual grounding from vision-language pre-training and open-vocabulary object detection. In Proceedings of the AAAI Conference on Artificial Intelligence, 2024.
@@
- [49] S. Srinidhi, E. Lu, and A. Rowe. XaiR: An XR platform that integrates large language models with the physical world. In IEEE International Symposium on Mixed and Augmented Reality (ISMAR), 2024.
- [50] D. Tochilkin, D. Pankratz, Z. Liu, Z. Huang, A. Letts, Y. Li, D. Liang, C. Laforte, V. Jampani, and Y. Cao. TripoSR: Fast 3D object reconstruction from a single image. arXiv preprint arXiv:2403.02151, 2024.
- [51] W. Tseng, E. Bonnail, M. McGill, M. Khamis, E. Lecolinet, S. Huron, and J. Gugenheimer. The dark side of perceptual manipulations in virtual reality. In Proceedings of the ACM CHI Conference on Human Factors in Computing Systems, 2022.
@@
- [57] J. Xiang, Z. Lv, S. Xu, Y. Deng, R. Wang, B. Zhang, D. Chen, X. Tong, and J. Yang. Structured 3D latents for scalable and versatile 3d generation. In Proceedings of the IEEE/CVF Computer Vision and Pattern Recognition Conference (CVPR), 2025.